OpenStudy (anonymous):

*MEDAL* Which equation best represents the regression line for the data given in the table below? x | y 1 | 13 2 | 17 3 | 14 4 | 8 5 | 3 6 | -2 7 | 2 (Choices): y = -3x - 20 y = -3x + 20 y = 3x + 20 y = 3x - 20

OpenStudy (austinl):

Do we know if this is indeed a line?

OpenStudy (anonymous):

is there a quick way of telling without having to graph it?

OpenStudy (austinl):

Well, I have a way in mind... but it will only work if this is a straight line, which I am assuming that it is.

OpenStudy (anonymous):

The choices kind of sum-up that it's a line :P

OpenStudy (austinl):

Ok, so my method is a little involved... but hey, you have time right? :P

OpenStudy (anonymous):

Yes, im good!

OpenStudy (austinl):

Wait, this is a data plot... where you would plot all the points and then find a line that fits it the best.

OpenStudy (anonymous):

Should I graph these then?

OpenStudy (austinl):

Well, plot the points.

OpenStudy (austinl):

There is a complex formula that I could give you... but I dunno if you'd want to use that.

OpenStudy (anonymous):

eeeh, well, if there's an easiler way of doing it...

OpenStudy (austinl):

Not really easier..... just different.

OpenStudy (anonymous):

Which ever one you think would work the best!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

\[ m=\frac{17-13}{2-1}=4 \]

OpenStudy (anonymous):

But it's better to do the first and last point. \[ m=\frac{2-13}{7-1}=-\frac{11}{6} \]

OpenStudy (anonymous):

\[(y=-3x+20)\]

OpenStudy (anonymous):

So then, the slope is 11/6? What do I use 11/6 for? :o

OpenStudy (anonymous):

the graph

OpenStudy (anonymous):

Oh I see! Thanks @wio @TravisCobb and @austinL

OpenStudy (austinl):

Sorry about that @NobodySpecial I got distracted :( Did you get to the answer?

OpenStudy (anonymous):

It's all good, @austinL :) yes, i got it.

OpenStudy (austinl):

Coolio!

OpenStudy (anonymous):

good

OpenStudy (anonymous):